(k) Find the value(s) of k such that the linear system X k has periodic solutions
Answer #1

-C A 6 X 2. Eigenvalues af 3-A - =o (3-a) (K-d ) - (2) (-6)- o P- 1. 2 d 3+ k)d +3 k + 12 = 0 d = (- (3k) ± J(-0+K))_ y (1 k+